OpenStudy (anonymous):

what about this one?? Use the binomial coefficient formula to calculate the binomial coefficient for the 8th term in a polynomial expansion when the power of the binomial is 10.

jimthompson5910 (jim_thompson5910):

\[\left(\begin{matrix}10 \\ 8\end{matrix}\right) =\frac{8!}{8!(10-8)!} = 45\]

OpenStudy (anonymous):

i dont get it the answer can not be 45

jimthompson5910 (jim_thompson5910):

This is using the combination formula. The value 45 is the coefficient. Oh and I made a typo, the 8! in the numerator should be 10!

OpenStudy (anonymous):

could you solve it with all the steps? please

jimthompson5910 (jim_thompson5910):

\[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10!}{8!(10-8)!}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10!}{8!2!}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10*9*8!}{8!2!}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10*9}{2!}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10*9}{2*1}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{90}{2}\] \[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=45\]

OpenStudy (anonymous):

720 22 120 240 those are the possibility

jimthompson5910 (jim_thompson5910):

Oh I made the silly mistake in thinking that x=8 is the 8th term, it's actually x=7 (since x starts at 0) So \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10!}{7!(10-7)!}=120\]

OpenStudy (anonymous):

ok could you solve this one with steps as well?

OpenStudy (anonymous):

What is the fourth term for the geometric sequence whose first term is 3 and whose ratio is 2?

jimthompson5910 (jim_thompson5910):

\[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10!}{7!(10-7)!}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10!}{7!3!}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10*9*8*7!}{7!3!}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10*9*8}{3!}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{10*9*8}{3*2*1}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=\frac{720}{6}\] \[\left(\begin{matrix}10 \\ 7\end{matrix}\right)=120\]

OpenStudy (anonymous):

What is the fourth term for the geometric sequence whose first term is 3 and whose ratio is 2?